HELLO! i am looking for help solving and understanding the process of getting to an answer with this problem.
You are studying the emerald ash borer and you want to determine the percentage of northern hardwood stands that have evidence of the emerald ash borer. How many stands must you sample in order to be 90% confident that your estimate is within 7 percentage points of the true population percentage? Research from 2 years ago gives a sample proportion of 63%.
Thank you in advance. I look forward to learning.

Answer by Boreal(15235)

the half-interval, or the margin of error, 0.07=z(0.95)*sqrt p(1-p)/n
in these circumstances 0.07=1.645*sqrt(0.63*0.37/n)
square both sides 0.0049=2.706*0.2331/n
n=2.706*0.2331/0.0049=128.73 round up to 129 stands